### v1

# When Trackers Date Fish: A Benchmark and Framework for Underwater Multiple Fish Tracking

Abstract
Multiple object tracking (MOT) technology has made significant progress in terrestrial applications, but underwater tracking scenarios remain underexplored despite their importance to marine ecology and aquaculture. We present Multiple Fish Tracking Dataset 2025 (MFT25), the first comprehensive dataset specifically designed for underwater multiple fish tracking, featuring 15 diverse video sequences with 408,578 meticulously annotated bounding boxes across 48,066 frames. Our dataset captures various underwater environments, fish species, and challenging conditions including occlusions, similar appearances, and erratic motion patterns. Additionally, we introduce Scale-aware and Unscented Tracker (SU-T), a specialized tracking framework featuring an Unscented Kalman Filter (UKF) optimized for non-linear fish swimming patterns and a novel Fish-Intersection-over-Union (FishIoU) matching that accounts for the unique morphological characteristics of aquatic species. Extensive experiments demonstrate that our SU-T baseline achieves state-of-the-art performance on MFT25, with 34.1 HOTA and 44.6 IDF1, while revealing fundamental differences between fish tracking and terrestrial object tracking scenarios. MFT25 establishes a robust foundation for advancing research in underwater tracking systems with important applications in marine biology, aquaculture monitoring, and ecological conservation. The dataset and codes are released at https://vranlee.github.io/SU-T/.
Figure 1: Challenges, features, and performance of the MFT25 dataset and SU-T baseline.
1 Introduction
Fish behavior monitoring and group dynamics analysis form essential technical foundations for marine ecological research, aquaculture optimization, and fishery resource management [8, 17]. With advancements in computer vision and deep learning, underwater multiple fish tracking (MFT) has emerged as a core technology for efficient, non-invasive observation [47, 23]. It enables quantitative analysis of fish movement patterns, group interactions, and environmental adaptation mechanisms by continuously tracking and associating individual targets across video sequences. MFT offers significant applications in endangered species protection, aquaculture density optimization, and marine ecosystem modeling [18, 41].
MFT is a specialized application of multiple object tracking (MOT), presents unique challenges in underwater environments [15, 10]. It aims to generate continuous trajectories of individual fish through reliable identification across video frames. Unlike single fish tracking, MFT must distinguish between numerous similar-looking fish and maintain consistent identity assignments despite rapid direction changes and frequent occlusions [4, 49]. The ability to resolve confusion between morphologically similar individuals in varying water conditions becomes critical to successful fish tracking, particularly in dense shoaling scenarios where individuals frequently cross paths [40].
Current methods largely rely on data-driven approaches, leveraging high-precision detectors to obtaizn real-time target positions [48, 21]. However, tracking fish in complex underwater environments presents several challenges, as shown in Fig. 1(a). On the one hand, high morphological similarity among individual fish combined with their erratic movement patterns frequently leads to identity switches and trajectory fragmentation [38, 22]. On the other hand, existing public datasets suffer from insufficient diversity and poor image quality [35, 34], limiting the development of tracking models with strong generalization capabilities across complex scenarios.
To address these challenges, we present Multiple Fish Tracking Dataset 2025 (MFT25), a large-scale dataset specifically designed for underwater multiple fish object tracking, alongside Scale-aware and Unscented Tracker (SU-T), an efficient, lightweight baseline model for online tracking. Our dataset and tracker aim to establish a robust foundation for advancing research in underwater object tracking systems with practical applications in marine ecology and aquaculture. Our main contributions are summarized as follows:
  • We introduce MFT25, the first comprehensive multiple fish tracking dataset featuring 15 diverse video sequences with 408,578 meticulously annotated bounding boxes across 48,066 frames, capturing various underwater environments, fish species, and challenging conditions including occlusions, rapid direction changes, and visually similar appearances.
  • We propose SU-T, a specialized tracking framework featuring an Unscented Kalman Filter (UKF) optimized for non-linear fish swimming patterns and a novel Fish-Intersection-over-Union (FishIoU) matching that accounts for the unique morphological characteristics and erratic movement behaviors of aquatic species.
  • We conduct extensive comparative experiments demonstrating that our tracker achieves state-of-the-art performance on MFT25, with 34.1 HOTA and 44.6 IDF1. Through quantitative analysis, we highlight the fundamental differences between fish tracking and land-based object tracking scenarios, as illustrated in Fig. 1(b) and (c).
2 MFT25: When Trackers Date Fish
2.1 Toward Fish Tracking
2.1.1 Related Methods
Fish tracking confronts distinctive challenges compared to terrestrial object tracking due to complex underwater environments, unique fish morphology, and erratic swimming behaviors [8]. Early approaches relied primarily on traditional computer vision techniques such as background subtraction [38] and object segmentation [17], which performed adequately only in controlled environments. Recent advances have shifted toward sophisticated deep learning implementations, including SiamRPN architectures [41], appearance-based models [23], graph-based tracking approaches [18], and advanced Swin Transformer architectures [47].
While MOT methods for terrestrial and underwater targets share fundamental principles, terrestrial tracking has been significantly more extensively developed [15]. Recent terrestrial MOT research has explored diverse technical aspects including multi-modal fusion strategies [20], adaptive frame rate processing [26, 12], computational efficiency optimizations [24], robust appearance modeling [37], depth information integration [27], and high-density scenario handling [19]. Contemporary MOT frameworks can be systematically categorized into three principal architectural approaches: Joint Detection and Embedding-based (JDE-based), Transformer-based, and Separate Detection and Embedding-based (SDE-based) methods.
JDE-based methods employ unified network architectures that simultaneously learn detection and tracking features, achieving training efficiency but often compromising feature specialization [42, 48, 21]. Transformer-based approaches leverage attention mechanisms to establish sophisticated spatiotemporal relationships, achieving high performance but introducing substantial computational overhead and memory requirements [11, 7, 44, 31, 22, 50]. SDE-based methods, predominantly extending the SORT framework [4], have gained widespread adoption by decoupling high-performance detectors from specialized appearance models, offering an optimal balance between accuracy and efficiency [49, 5, 43, 13, 1, 45].
Recent innovations in MOT have also explored camera pose estimation [46], generative diffusion models [29], and pixel-wise trajectory propagation [51]. However, these sophisticated techniques often demonstrate limited adaptability to underwater environments and insufficient real-time performance, rendering them suboptimal for practical fish tracking applications. The unique characteristics of underwater environments—including variable lighting conditions, water turbidity, and the distinctive motion patterns of aquatic species—necessitate specialized tracking solutions.
Figure 2: The framework of SU-T. The pipeline consists of three main components: a detector, an association module, and an optional Re-ID module. The detector generates bounding boxes and confidence scores for each frame. The optional Re-ID module extracts feature embeddings to enhance tracking accuracy. The association module uses the FishIoU to calculate matching costs between detected boxes and predicted boxes from the UKF.
2.1.2 Related Datasets
MOT datasets require frame-by-frame bounding box annotations with consistent identity information across extended video sequences, representing a significant annotation challenge [9, 32, 25]. Current MOT research predominantly focuses on terrestrial domains, resulting in well-established benchmarks for humans, vehicles, and animals, such as the MOT challenge series [10], DanceTrack [40], BEE24 [6], and CTMC [2].
While several fish-oriented video datasets exist, including Fish4Knowledge, DeepFish, and SeaCLEF [8], they present significant limitations for modern tracking applications. These early datasets typically suffer from low resolution, poor visibility conditions that obscure fish identities, and inconsistent annotation formats that impede effective model training. More recent standardized fish tracking datasets such as BrackishMOT [35], 3D-ZeF [34], and MFT22 [22] have emerged with consistent annotation protocols. However, these datasets remain limited in both environmental diversity and scale, typically featuring simplified scenarios with controlled conditions. The absence of comprehensive, high-quality, and standardized fish tracking datasets thus represents a critical bottleneck constraining significant advances in underwater MFT research.
2.2 Dataset Construction
The MFT25 dataset was captured using imaging equipment of Canon $EOSR6$ and Sony $\alpha 7M3$, across diverse aquaculture environments. Recording locations encompassed both industrial circulating water aquaculture ponds and controlled laboratory tanks to ensure environmental diversity. The dataset features multiple fish species with distinctly different morphologies, including commercially valuable groupers and ornamental koi at various developmental stages, providing substantial variation in appearance characteristics.
To ensure comprehensive scenario coverage, we systematically deployed multiple camera configurations, including both overhead and horizontal perspectives, across varied illumination conditions from daylight to nocturnal settings. The dataset consists exclusively of authentic footage without synthetic augmentation, preserving the natural complexity of scenarios. All bounding box annotations were created using DarkLabel [^1] https://github.com/darkpgmr/DarkLabel software through manual selection and verification processes. The resulting MFT25 dataset encompasses 15 diverse video sequences containing 223 distinct fish trajectories across 48,066 frames, with a total of 408,578 precisely annotated bounding boxes. This represents a substantial advancement in scale, containing 2.6-9.1 times more annotated instances compared to previous fish tracking datasets. Table 1 presents a comprehensive statistical comparison with existing fish tracking benchmarks.
Table 1: Quantitative comparison of MFT datasets.
Dataset BrackishMOT 3D-ZeF MFT22 MFT25(Ours)
Clips 86 8 10 15
Tracks 600 32 234 223
FPS 25 60 25 25
Frames 12,466 28,800 9,100 48,066
Boxes 44,809 86,400 155,437 408,578
Moreover, we analyze the distinctive movement characteristics of fish in our dataset through detailed quantitative analysis in Section 4.4.
3 SU-T: A MFT Baseline
3.1 Framework
To address the unique challenges of MFT, we propose Scale-aware and Unscented Tracker (SU-T), a specialized baseline following the SDE paradigm. As illustrated in Fig. 2, our framework comprises three primary components: a detector, an association module, and an optional Re-Identification (Re-ID) module. The processing pipeline begins with video frames being fed into the detector, which generates bounding boxes with corresponding confidence scores. These detections are then processed by the association module, where our specialized FishIoU metric calculates matching costs between detected boxes and predictions from the UKF. The Hungarian algorithm performs optimal assignment to update existing trajectories and establish new tracks when necessary. To address the challenge of visually similar fish, SU-T integrates an optional Re-ID module that extracts discriminative feature embeddings. These embeddings work synergistically with the FishIoU metric during association, significantly enhancing tracking accuracy by maintaining consistent identities even when fish exhibit nearly identical appearances.
3.2 Unscented Kalman Filter
Figure 3: Unscented Kalman Filter (UKF) motion model used in SU-T for predicting fish movement in complex underwater environments.
The Unscented Kalman Filter (UKF) is particularly well-suited for tracking fish due to their non-linear motion patterns, as shown in Fig. 3. Unlike standard Kalman Filter, UKF uses a deterministic sampling technique to handle non-linearities. The three core mathematical components of our UKF implementation are as follows.
3.2.1 Sigma Points Generation
For state vector $\mathbf{x}\in\mathbb{R}^{n}$ with covariance $\mathbf{P}$, we generate $12n+$ 1 sigma points:
$\mathcal{X}_{0}=\mathbf{x}$ (1)
$\mathcal{X}_{i}=\mathbf{x}+\left(\sqrt{(n+\lambda)\mathbf{P}}\right)_{i},\quad i=1,\ldots,n$ (2)
$\mathcal{X}_{i+n}=\mathbf{x}-\left(\sqrt{(n+\lambda)\mathbf{P}}\right)_{i},\quad i=1,\ldots,n$ (3)
where $\lambda=\alpha^{2}(n+\kappa)-n$ is a scaling parameter, $\alpha$ controls spread of points, $\kappa$ is a secondary parameter (typically 3-n), and $\left(\sqrt{(n+\lambda)\mathbf{P}}\right)_{i}$ is the $i$-th column of the matrix square root.
3.2.2 Prediction Step
Each sigma point is propagated through the non-linear state transition function $\mathbf{f}$:
$\mathcal{X}_{k|k-1}^{i}=\mathbf{f}(\mathcal{X}_{k-1}^{i}),\quad i=0,\ldots,2n$ (4)
$\hat{\mathbf{x}}_{k|k-1}=\sum_{i=0}^{2n}w_{m}^{i}\mathcal{X}_{k|k-1}^{i}$ (5)
$\mathbf{P}_{k|k-1}=\sum_{i=0}^{2n}w_{c}^{i}[\mathcal{X}_{k|k-1}^{i}-\hat{\mathbf{x}}_{k|k-1}][\mathcal{X}_{k|k-1}^{i}-\hat{\mathbf{x}}_{k|k-1}]^{T}+\mathbf{Q}_{k}$ (6)
where $w_{m}^{i}$ and $w_{c}^{i}$ are weight coefficients for mean and covariance, and $\mathbf{Q}_{k}$ is the process noise covariance.
3.2.3 Measurement Update Step
The predicted sigma points are transformed through the measurement function $\mathbf{h}$:
$\mathcal{Z}_{k|k-1}^{i}=\mathbf{h}(\mathcal{X}_{k|k-1}^{i}),\quad i=0,\ldots,2n$ (7)
$\hat{\mathbf{z}}_{k|k-1}=\sum_{i=0}^{2n}w_{m}^{i}\mathcal{Z}_{k|k-1}^{i}$ (8)
$\mathbf{P}_{zz}=\sum_{i=0}^{2n}w_{c}^{i}[\mathcal{Z}_{k|k-1}^{i}-\hat{\mathbf{z}}_{k|k-1}][\mathcal{Z}_{k|k-1}^{i}-\hat{\mathbf{z}}_{k|k-1}]^{T}+\mathbf{R}_{k}$ (9)
$\mathbf{P}_{xz}=\sum_{i=0}^{2n}w_{c}^{i}[\mathcal{X}_{k|k-1}^{i}-\hat{\mathbf{x}}_{k|k-1}][\mathcal{Z}_{k|k-1}^{i}-\hat{\mathbf{z}}_{k|k-1}]^{T}$ (10)
$\mathbf{K}_{k}=\mathbf{P}_{xz}\mathbf{P}_{zz}^{-1}$ (11)
$\hat{\mathbf{x}}_{k}=\hat{\mathbf{x}}_{k|k-1}+\mathbf{K}_{k}(\mathbf{z}_{k}-\hat{\mathbf{z}}_{k|k-1})$ (12)
$\mathbf{P}_{k}=\mathbf{P}_{k|k-1}-\mathbf{K}_{k}\mathbf{P}_{zz}\mathbf{K}_{k}^{T}$ (13)
where $\mathbf{z}_{k}$ is the actual measurement, $\mathbf{R}_{k}$ is the measurement noise covariance, and $\mathbf{K}_{k}$ is the Kalman gain.
3.3 FishIoU: Scale-aware Association
The unique morphology and movement patterns of fish present significant challenges for standard object association. We introduce Fish-Intersection-over-Union (FishIoU), a specialized association IoU that accounts for the elongated body structure, erratic motion patterns, and size variations common in fish species.
Given two bounding boxes $B_{1}=[x_{1},y_{1},x_{2},y_{2}]$ and $B_{2}=[x^{\prime}_{1},y^{\prime}_{1},x^{\prime}_{2},y^{\prime}_{2}]$, we first compute the standard IoU as:
$\text{IoU}=\frac{|B_{1}\cap B_{2}|}{|B_{1}\cup B_{2}|}$ (14)
where $|B_{1}\cap B_{2}|$ represents the intersection area and $|B_{1}\cup B_{2}|$ the union area. Then, to account for fish morphology, we incorporate a center distance penalty:
$d_{c}=\frac{(c_{x}-c^{\prime}_{x})^{2}+(c_{y}-c^{\prime}_{y})^{2}}{d_{\text{diag}}^{2}}$ (15)
where $(c_{x},c_{y})$ and $(c^{\prime}_{x},c^{\prime}_{y})$ are the centers of $B_{1}$ and $B_{2}$ respectively, and $d_{\text{diag}}^{2}$ is the squared diagonal length of the enclosing box. Considering fish bodies typically have an elongated structure with important features concentrated in the front, we define a central region for each box with asymmetric insets to emphasize this characteristic:
$B_{1}^{c}$ $=[x_{1}+\alpha w_{1},y_{1}+\beta h_{1},x_{2}-\gamma w_{1},y_{2}-\beta h_{1}]$ (16)
$B_{2}^{c}$ $=[x^{\prime}_{1}+\alpha w_{2},y^{\prime}_{1}+\beta h_{2},x^{\prime}_{2}-\gamma w_{2},y^{\prime}_{2}-\beta h_{2}]$ (17)
where $w_{1},h_{1}$ and $w_{2},h_{2}$ are the width and height of $B_{1}$ and $B_{2}$ respectively, and $\alpha,\beta,\gamma$ are constant factors determined empirically based on fish morphological characteristics. By default, $\alpha=0.15$, $\beta=0.3$, and $\gamma=0.25$. The central IoU is then calculated as:
$\text{cIoU}=\frac{|B_{1}^{c}\cap B_{2}^{c}|}{|B_{1}^{c}\cup B_{2}^{c}|}$ (18)
To account for consistent fish orientation, we consider the aspect ratio consistency:
$\alpha_{r}=\frac{\min(r_{1},r_{2})}{\max(r_{1},r_{2})}$ (19)
where $r_{1}=\frac{w_{1}}{h_{1}}$ and $r_{2}=\frac{w_{2}}{h_{2}}$ are the aspect ratios of the two boxes.
Additionally, we incorporate area ratio consistency, since the size of the fish does not change abruptly between frames:
$\alpha_{a}=\frac{\min(a_{1},a_{2})}{\max(a_{1},a_{2})}$ (20)
where $a_{1}=w_{1}\times h_{1}$ and $a_{2}=w_{2}\times h_{2}$ are the areas of the two boxes.
For small targets, we apply a scale factor to reduce the center distance penalty:
$s=1-e^{-\frac{\min(a_{1},a_{2})}{1000}}$ (21)
The final FishIoU metric combines these components with specific weights optimized for fish tracking:
$\text{FishIoU}=\omega_{1}\cdot\text{IoU}+\omega_{2}\cdot\text{cIoU}+\omega_{3}\cdot\alpha_{r}+\omega_{4}\cdot\alpha_{a}-\omega_{5}\cdot s\cdot d_{c}$ (22)
where $\omega_{1}=1.0$, $\omega_{2}=0.3$, $\omega_{3}=0.1$, $\omega_{4}=0.2$, and $\omega_{5}=0.4$ are weights determined empirically through extensive experiments.
3.4 Assocation
Our association strategy employs progressive confidence-based processing that significantly reduces identity switches while maintaining computational efficiency. Algorithm 1 presents the complete multi-level cascade tracking process, which integrates our UKF motion prediction and FishIoU matching with a cascaded association strategy for aquatic environments.
Building upon frameworks from HybridSORT [45], we adopt dual-confidence matching of ByteTrack [49] and heuristic observation-centric recovery of OC-SORT [5], and implements three association stages. In the first stage, high-confidence detections are matched with existing tracks using our specialized FishIoU, establishing reliable primary associations even when fish exhibit rapid direction changes. The second stage associates remaining tracks with low-confidence detections, effectively recovering temporarily occluded targets while filtering false positives induced by water turbidity and reflections. The final stage attempts to reconnect tracks with their historical appearances, addressing the frequent, abrupt directional changes and non-linear swimming patterns characteristic of fish locomotion.
For cost calculation, our framework integrates both spatial and appearance information when available. Spatially, the FishIoU outperforms standard IoU by incorporating fish-specific morphological features into the matching process. When enabled, the Re-ID module provides discriminative appearance embeddings that effectively differentiate between visually similar individuals swimming in close proximity.
Input: Detections $\mathcal{D}$ with scores, Existing tracks $\mathcal{T}$
Output: Updated tracks $\mathcal{T}$
/* Prediction */
for $T_{j}\in\mathcal{T}$ do
       $\hat{B}_{j},s_{j}\leftarrow$ UKF.predict($T_{j}$) /* Predict box and score */
end for
/* First Association: High-confidence */
$\mathcal{D}_{high}\leftarrow\{d_{i}\in\mathcal{D}:score(d_{i})>\tau_{high}\}$
$\mathbf{C}\leftarrow$ FishIoU($\mathcal{D}_{high}$, $\{\hat{B}_{j}\}$) /* Base cost */
if Use Re-ID then
      $\mathbf{C}\leftarrow\omega_{1}\mathbf{C}+\omega_{2}$ EmbeddingDistance($\mathcal{D}_{high}$, $\mathcal{T}$)
end if
$\mathcal{M}_{1},\mathcal{U}_{\mathcal{D}},\mathcal{U}_{\mathcal{T}}\leftarrow$ Hungarian($-\mathbf{C}$)
for $(i,j)\in\mathcal{M}_{1}$ do
      $T_{j}$.update($d_{i}$)
end for
/* Second Association: Low-confidence */
$\mathcal{D}_{low}\leftarrow\{d_{i}\in\mathcal{D}:\tau_{low}<score(d_{i})<\tau_{high}\}$
$\mathbf{C}_{iou}\leftarrow$ FishIoU($\mathcal{D}_{low}$, $\{\hat{B}_{j}:T_{j}\in\mathcal{U}_{\mathcal{T}}\}$)
if Use Re-ID then
      $\mathbf{C}_{iou}\leftarrow\mathbf{C}_{iou}+\lambda\cdot$ EmbeddingDistance($\mathcal{D}_{low}$, $\{T_{j}\in\mathcal{U}_{\mathcal{T}}\}$)
end if
$\mathcal{M}_{2}\leftarrow$ Hungarian($-\mathbf{C}_{iou}$)
for $(i,j)\in\mathcal{M}_{2}$ where $\mathbf{C}_{iou}[i,j]>\tau_{iou}$ do
       $T_{j}$.update($\mathcal{D}_{low}[i]$) /* Update state */
       $\mathcal{U}_{\mathcal{T}}\leftarrow\mathcal{U}_{\mathcal{T}}\setminus\{T_{j}\}$
end for
/* Third Association: Last-chance */
$\mathbf{C}_{last}\leftarrow$ FishIoU($\mathcal{D}_{high}[\mathcal{U}_{\mathcal{D}}]$, $\{T_{j}.last\_observation:T_{j}\in\mathcal{U}_{\mathcal{T}}\}$)
for $(i,j)\in$ Hungarian($-\mathbf{C}_{last}$) where $\mathbf{C}_{last}[i,j]>\tau_{iou}$ do
       $T_{j}$.update($\mathcal{D}_{high}[\mathcal{U}_{\mathcal{D}}[i]]$)
       Remove $i$ from $\mathcal{U}_{\mathcal{D}}$, $T_{j}$ from $\mathcal{U}_{\mathcal{T}}$
end for
/* Finalize */
Update all $T_{j}\in\mathcal{U}_{\mathcal{T}}$ without observation
Initialize new tracks from $\mathcal{D}_{high}[\mathcal{U}_{\mathcal{D}}]$
Remove tracks with $time\_since\_update>max\_age$
return $\mathcal{T}$
Algorithm 1 Multi-Level Cascade Tracking
Table 2: Comparison of different tracking methods on the MFT25 dataset. indicates the integration of Re-ID module. The best two results are bolded and underlined respectively. Same as follows.
Method Class HOTA↑ IDF1↑ MOTA↑ AssA↑ DetA↑ IDs↓ IDFP↓ IDFN↓ Frag↓
FairMOT [48] JDE 22.226 26.867 47.509 13.910 35.606 939 58198 113393 3768
CMFTNet [21] JDE 22.432 27.659 46.365 14.278 35.452 1301 64754 111263 2769
Deep-OC-SORT [30] SDE 24.848 34.176 46.721 17.537 35.373 550 53478 104024 3659
OC-SORT [5] SDE 25.017 34.620 46.706 17.783 35.369 550 52934 103495 3651
TFMFT [22] Transformer 25.440 33.950 49.725 17.112 38.059 719 63125 102378 3251
BoT-SORT [1] SDE 26.848 36.847 49.108 19.446 37.241 500 57581 99181 2704
TransCenter [44] Transformer 27.896 30.278 68.693 30.255 30.301 807 101223 101002 1992
SORT [4] SDE 29.063 34.119 69.038 16.952 50.195 778 88928 96815 1726
TrackFormer [31] Transformer 30.361 35.285 74.609 17.661 52.649 718 89391 94720 1729
TransTrack [39] Transformer 30.426 35.215 68.983 18.525 50.458 1116 96045 93418 2588
ByteTrack [49] SDE 31.758 40.355 69.586 20.392 49.712 489 80765 87866 1555
HybridSORT [45] SDE 32.258 38.421 68.905 20.936 49.992 613 85924 90022 1931
HybridSORT [45] SDE 32.705 41.727 69.167 21.701 49.697 562 79189 85830 1963
SU-T (Ours) SDE 33.351 41.717 68.450 22.425 49.943 607 83111 84814 2006
SU-T (Ours) SDE 34.067 44.643 68.958 23.594 49.531 544 76440 81304 2011
4 Experiments
4.1 Implementation Details
For all experiments, we employed the same YOLOX-X [14] detector and the same Re-ID network [16] for SDE-based models, trained with consistent hyperparameter configurations, following the established protocols from ByteTrack [49] and BoT-SORT [1]. All models were trained on the MFT25 training set with a Tesla A100 GPU, with performance on the test set evaluated using standard MOT metrics, including the comprehensive HOTA [28] metric alongside traditional CLEAR [3] metrics such as MOTA, IDF1, and ID switches (IDs).
4.2 Benchmark Results
We compare our proposed SU-T baseline with state-of-the-art MOT and MFT methods on the MFT25 dataset. Table 2 presents the comprehensive comparison results. Our method achieves the best overall performance with a HOTA score of 34.1, surpassing the previous best method by 1.4. This improvement demonstrates the effectiveness of our specialized fish tracking approach compared to general-purpose tracking methods. The superiority of our method is particularly evident in association metrics, where SU-T achieves the highest IDF1 score of 44.6 with Re-ID module, significantly outperforming other methods. This indicates that our approach is more effective at maintaining consistent fish identities across frames, which is crucial for accurate trajectory analysis in complex underwater scenarios.
Although Transformer-based TrackFormer [31] achieving the highest MOTA score of 74.6. However, it exhibit relatively weaker performance in identity preservation metrics such as IDF1 and AssA. Besides, transformer-based trackers impose substantial computational overhead, rendering them impractical for real-time underwater monitoring applications. The performance gap between conventional terrestrial-focused trackers and SU-T validates our hypothesis that underwater tracking scenarios necessitate domain-specific adaptations. Visual comparisons of various tracking methods on MFT25 are illustrated in Fig. 4.
In addition, we conducted additional experiments to evaluate the generalization capability of SU-T on mainstream land-based tracking benchmarks MOT17 and MOT20 [10]. Our baseline achieved 60.4 and 56.5 HOTA, respectively. [^2] The results have been registered at https://motchallenge.net/
Figure 4: Tracking performance of various trackers on the MFT25 dataset. F_NUM denotes false tracked number, including IDFN, IDFP, and IDs. Best viewed in color.
Figure 5: Visualization of SU-T tracking performance across diverse scenarios in the MFT25 dataset. Each tracked fish maintains consistent identity despite challenging underwater conditions including variable visibility, population density, and complex backgrounds. Best viewed in color.
4.3 Ablation Studies
Table 3: Results on different Re-ID models with standard Kalman Filter and IoU association.
Method IBN HOTA↑ IDF1↑ MOTA↑ AssA↑ IDs↓
SBS-R50 30.950 39.937 68.780 19.599 713
SBS-R50 30.560 37.919 68.909 19.010 659
SBS-R101 30.270 38.104 68.796 18.599 678
SBS-R101 30.684 39.996 68.912 19.134 638
SBS-S50 32.705 41.727 69.167 21.701 562
SBS-S50 32.412 40.977 69.030 21.183 558
SBS-S101 33.842 43.748 69.043 23.154 550
SBS-S101 31.900 40.201 69.212 20.610 584
We conducted extensive ablation experiments to evaluate the components. Table 3 shows the impact of different Re-ID model on tracking performance. The SBS-S101 [16] model achieves the best overall performance with 33.8 HOTA. Notably, the Instance-Batch Normalization (IBN) [33] variants do not consistently improve performance, suggesting that domain adaptation techniques optimized for terrestrial tracking may not transfer effectively to underwater scenarios. As a result, we adopt SBS-S101 as our Re-ID module.
Table 4 compares different IoU for association cost calculation. Our specialized FishIoU outperforms other association methods, achieving the highest 33.4 HOTA and 41.7 IDF1 scores. When combined with Re-ID module, performance further improves across all metrics, confirming that our FishIoU better handles the unique morphology and movement patterns of fish species.
Table 4: Ablation study comparing different IoU for association cost calculation. Center and IoU represent the center points distance and the standard IoU, respectively.
Method HOTA↑ IDF1↑ MOTA↑ AssA↑ IDs↓ IDFP↓ IDFN↓
Center 28.865 37.348 66.313 17.410 1273 88585 91297
IoU 32.790 40.098 68.839 21.573 579 84648 87347
CIoU [53] 30.720 39.598 67.425 19.325 727 87422 87556
DIoU [52] 30.764 39.575 67.519 19.326 728 87344 87617
HMIoU [45] 32.258 38.421 68.905 20.936 613 85924 90022
GIoU [36] 32.885 39.957 68.798 21.686 573 84896 87538
FishIoU 33.351 41.717 68.450 22.425 607 83111 84814
FishIoU 33.581 43.268 68.989 22.779 547 78473 83243
Table 5 investigates the effectiveness of different motion models combined with our association metrics. The UKF consistently outperforms other motion models including standard Kalman Filter (KF), Adaptive Kalman Filter (AKF), and Strong Tracking Filter (STF) across both HMIoU [45] and FishIoU association. This validates our hypothesis that non-linear motion models are more appropriate for capturing the complex swimming patterns of fish. The best performance is achieved by combining UKF with FishIoU and Re-ID, resulting in a HOTA of 34.1 and IDF1 of 44.6.
Table 6 presents the performance of our SU-T baseline across individual sequences in the MFT25 dataset, revealing substantial variation in tracking difficulty. The PF-002 sequence demonstrates exceptional performance with 65.2 HOTA, 98.7 IDF1, and none IDs, highlighting the effectiveness of SU-T under favorable conditions with distinct fish appearances and predictable movements. Similarly, BT-004 achieves strong results with 51.2 HOTA and 68.9 IDF1, demonstrating robust tracking capabilities in moderately challenging scenarios where fish maintain relatively consistent swimming patterns.
In contrast, sequences like SN-011 exhibit significantly lower performance with 20.3 HOTA, due to challenging conditions including numerous occlusions and highly erratic movements. MSK-003 shows moderate 34.2 HOTA but suffers from higher 134 IDs, indicating difficulties in maintaining consistent identities in crowded environments. SN-009 and SN-011 both demonstrate substantial identification failures, with high IDFN values of 30,853 and 23,575, respectively. This reveals the limitations of current tracking approaches in visibility-constrained underwater environments.
To visualize the practical effectiveness of SU-T, Fig. 5 presents tracking results across diverse scenarios in the MFT25 dataset. The visualizations demonstrate that SU-T maintains consistent identity assignments despite varying challenges including different fish densities, species morphologies, and environmental conditions.
Table 5: Ablation study on different motion models and association IoUs.
Method HMIoU FishIoU HOTA↑ IDF1↑ MOTA↑ AssA↑ IDs↓
KF 32.258 38.421 68.905 20.936 613
AKF 28.769 33.689 67.827 16.954 1031
STF 31.105 36.911 69.161 19.445 667
UKF 32.406 38.408 68.933 21.096 609
UKF 33.737 43.880 69.057 23.063 528
KF 33.051 41.041 68.503 22.022 612
AKF 22.551 24.017 65.535 10.682 2368
STF 31.601 38.137 68.694 20.153 663
UKF 33.201 41.644 68.451 22.261 609
UKF 34.067 44.643 68.958 23.594 544
Table 6: Performance of our baseline method on individual sequences in the MFT25 dataset.
Clip HOTA↑ IDF1↑ MOTA↑ AssA↑ IDs↓ IDFP↓ IDFN↓
BT-002 35.961 58.562 55.923 28.788 35 1917 2095
BT-004 51.244 68.933 93.028 42.125 21 2299 2408
MSK-003 34.188 53.067 50.393 28.930 134 12488 15001
PF-002 65.162 98.650 97.300 65.399 0 82 80
SN-009 27.169 28.799 82.766 12.692 173 29339 30853
SN-011 20.264 26.165 47.423 10.421 161 27146 23575
SN-014 40.046 47.718 84.339 26.700 83 9840 10802
Total 33.351 41.717 68.450 22.425 607 83111 84814
Figure 6: Temporal evolution of angular velocities across different tracking datasets. MFT25 demonstrates significantly higher directional variability and more erratic motion patterns compared to other tracking scenarios.
Figure 7: Temporal evolution of average object speeds across different tracking datasets. MFT25 demonstrates consistently higher velocities and greater variability compared to other tracking scenarios.
4.4 Statistical Analysis
We conducted comparative statistical analyses across diverse tracking datasets, including pedestrians [10], dancers [40], bees [6], and cells [2], to quantify the unique motion characteristics that distinguish fish tracking from other domains. Fig. 1(c) illustrates the distribution of motion directions, while Fig. 6 presents the angular velocity profiles across datasets. Our analysis reveals statistically significant differences in movement patterns. The MOT17 [10] dataset demonstrates remarkably stable angular velocities and predominantly linear motion patterns typical of pedestrian movement. In contrast, MFT25 exhibit substantially higher angular velocity variations, with more chaotic directional changes while maintaining consistently high swimming speeds. This erratic behavior stems from the hydrodynamic properties of underwater locomotion and environmental factors, creating inherently non-linear motion trajectories.
As shown in Fig. 7, MFT25 features comparatively higher average motion speeds, while the CTMC [2] dataset displays the lowest velocity profiles. These fundamental kinematic differences underscore why conventional tracking algorithms optimized for terrestrial scenarios often underperform in underwater environments, highlighting the need for specialized fish tracking approaches that can accommodate these unique movement characteristics.
5 Conclusion
In this paper, we introduces a unified underwater MFT benchmark and a specialized tracking framework for fish morphology and erratic swimming patterns. Our baseline achieves state-of-the-art performance with 34.1 HOTA, significantly outperforming other trackers. Statistical analysis reveals fundamental differences between fish and terrestrial tracking scenarios, highlighting the necessity for specialized underwater approaches. However, significant challenges remain in handling visually similar fish appearances, extreme density scenarios, and highly erratic swimming patterns, as evidenced by performance variations across sequences. Nevertheless, MFT25 and SU-T establish a robust foundation for advancing underwater tracking systems with important applications in marine biology, aquaculture monitoring, and ecological conservation.